# Probabilistic-based Feature Embedding of 4-D Light Fields for Compressive Imaging and Denoising

### v1

@@ 6.3.3 Comparison of Running Time
We compared the inference time (in seconds) of different methods for LF denoising under HCI+Inria dataset, and Table 6 lists the results. All methods were tested on a desktop with Intel CPU i7-8700 @ 3.70GHz, 32 GB RAM and NVIDIA GeForce RTX 3090. As shown in Table 6, Ours is faster than Alain and Smolic (2017), Tomita et al. (2022), Chen et al. (2018a) and Guo et al. (2021) but slower than Ren et al. (2021).
@@ 6.3.4 Ablation Study
We conducted ablation studies on the HCI+Inria dataset with $\sigma=20$ to verify the effectiveness of the extended framework and the impact of number of stages and feature embedding units.
The number of stages and feature embedding units. We varied the number of iterative stages in the range from 1 to 4 with 6 fixed feature embedding units, and varied the number of feature embedding units, i.e., 4, 6, and 8 with 2 fixed iterative stages. As shown in Table 7, the denoising performance improves with the increase of the number of iterative stages. Moreover, the improvement is more obvious from 1 stage to 3 stages, while slight from 3 stages to 4 stages. What’s more, the performance improves with the number of feature embedding units increasing. To balance the denoising performance and the amount of network parameters, we finally choose 2 stages with 6 feature embedding units to process all the LF denoising tasks.
@@
Furthermore, to evaluate the effectiveness of the proposed probabilistic-based feature embedding, we replaced the probabilistic-based feature embedding $\mathcal{D}(\cdot)$ in each stage with either a stack of 3-D convolution layers or SAS layers while maintaining a similar model size to our framework. The resulting networks were denoted as Ours_Conv3D and Ours_SAS, respectively, and their performances were compared with Ours in Table 8.
Moreover, we proposed a baseline framework without the proposed modules and replaced the feature embedding units with SAS. As shown in Table 8, it can be observed that without the proposed noise suppression module, the value of PSNR decreased by approximately $0.45dB$, thereby validating its effectiveness. Furthermore, Ours achieved the best performance compared with 3-D convolution and SAS, demonstrating the advantage of the proposed probabilistic-based feature embedding. Moreover, the effectiveness of the overall framework was validated by comparing the baseline with Ours, where it can be seen that the value of PSNR improved by $2.3dB$.
@@
Table 8: Effectiveness of the proposed framework in terms of average PSNR, SSIM.
@@
Ours_Conv3D(1) 36.82 0.920 2.02
Ours_Conv3D(2) 36.54 0.922 2.68
Ours_SAS 38.82 0.950 2.32
Ours 40.50 0.962 2.39
@@ 7 Conclusion
We have presented a novel adaptive probabilistic-based feature embedding for LFs. To verify its effectiveness, we evaluated its performance on compressive LF reconstruction and LF denoising, respectively. For compressive LF imaging, we proposed a physically-interpretable cycle-consistent framework that incorporates probabilistic-based feature embedding. Our approach can reconstruct 4-D LFs from 2-D measurements with higher quality, improving the PSNR value by approximately 4.5 dB while preserving the LF parallax structure better than state-of-the-art methods. Furthermore, we demonstrate the efficacy of our probabilistic-based feature embedding approach for LF denoising through a carefully designed network. Extensive experiments on diverse datasets demonstrate that our method outperforms state-of-the-art approaches by up to $1.1$ dB.

### v3

@@ 6.3.4 Comparison of Running Time
We compared the inference time (in seconds) of different methods for LF denoising under HCI+Inria dataset, and Table 6 lists the results. All methods were tested on a desktop with Intel CPU i7-8700 @ 3.70GHz, 32 GB RAM and NVIDIA GeForce RTX 3090. As shown in Table 6, Ours is faster than Alain and Smolic (2017), Tomita et al. (2022), Chen et al. (2018a) and Guo et al. (2022) but slower than Ren et al. (2021) and Wang et al. (2023c).
@@ 6.3.5 Ablation Study
We conducted ablation studies on the HCI+Inria dataset with $\sigma=20$ to verify the effectiveness of the extended framework and the impact of the number of stages and feature embedding units.
The number of stages and feature embedding units. We varied the number of iterative stages in the range from 1 to 4 with 6 fixed feature embedding units, and varied the number of feature embedding units, i.e., 4, 6, and 8 with 2 fixed iterative stages. As shown in Table 7, the denoising performance improves with the increase in the number of iterative stages. Moreover, the improvement is more obvious from 1 stage to 3 stages, while slight from 3 stages to 4 stages. What’s more, the performance improves with the number of feature embedding units increasing. To balance the denoising performance and the number of network parameters, we finally choose 2 stages with 6 feature embedding units to process all the LF denoising tasks.
The application of the proposed probabilistic-based feature embedding module to various LF tasks requires a balance between performance and computational efficiency. As a general guideline, we recommend maintaining the number of network parameters within the range of $2M\sim 4M$. This can be achieved using configurations such as $T=3,J=6,C=64$ or $T=6,J=6,C=32$, where $C$ represents the number of feature channels. These initial configurations can be fine-tuned further as per specific requirements.
@@
Furthermore, to evaluate the effectiveness of the proposed probabilistic-based feature embedding, we replaced the probabilistic-based feature embedding $\mathcal{D}(\cdot)$ in each stage with either a stack of 3-D convolution layers or SAS layers while maintaining a similar model size to our framework. The resulting networks were denoted as DN-Conv3D and DN-SAS, respectively, and their performances were compared with the proposed DN-Net in Table 9.
Table 8: Comparative analysis of the effectiveness of the proposed PFE module on LF spatial and angular super-resolution (SR). We selected 2×2$\times$2 × spatial SR with $5\times 5$ SAIs and $2\times 2\rightarrow 7\times 7$ angular SR for the comparison. The table presents the PSNR/SSIM results for two tasks, DistgSSR and DistgASR, with and without the integration of PFE.
2×2$\times$2 × Model EPLF HCInew HCIold Inria STFgtr
Spatial SR DistgSSR 34.58/0.977 37.76/0.979 44.77/0.995 36.48/0.985 40.29/0.994
DistgSSR-PFE 34.85/0.978 37.92/0.979 44.86/0.995 36.69/0.986 40.55/0.994
$2\times 2\rightarrow 7\times 7$ Model HCInew HCIold 30scenes Occlusion Reflective
Angular SR DistgASR 34.28/0.970 42.12/0.977 43.36/0.995 39.15/0.990 38.85/0.977
DistgASR-PFE 34.52/0.972 42.33/0.979 43.64/0.995 39.40/0.991 39.08/0.978
Moreover, we proposed a baseline framework without the proposed modules and replaced the feature embedding units with SAS. As shown in Table 9, it can be observed that without the proposed noise suppression module, the value of PSNR decreased by approximately $0.45dB$, thereby validating its effectiveness. Furthermore, the proposed DN-Net outperformed the 3-D convolution, SAS, and template network (DN-Template) in our evaluations, demonstrating the advantage of the proposed probabilistic-based feature embedding. Moreover, the effectiveness of the overall framework was validated by comparing the baseline with the proposed DN-Net, where it can be seen that the value of PSNR improved by $2.3dB$.
@@
Table 9: Effectiveness of the proposed framework in terms of average PSNR and SSIM values.
@@
DN-Conv3D(1) 36.82 0.920 2.02
DN-Conv3D(2) 36.54 0.922 2.68
DN-SAS 38.82 0.950 2.32
DN-Template 40.27 0.960 2.41
DN-Net 40.50 0.962 2.39
6.4 Evaluation on 4-D LF Super-Resolution
In addition to the coded aperture-based compressive LF reconstruction, LF super-resolution approaches, including spatial and angular super-resolution techniques, present an alternative route to achieving high-quality LFs. The spatial-angular trade-offs present in LF cameras have incited the development of several LF super-resolution methods to enhance spatial or angular resolution.
To further validate the efficacy of our Probabilistic-based Feature Embedding (PFE) approach, we conducted experiments based on a recent work Wang et al. (2023a). We specifically applied PFE to two distinct tasks, namely DistgSSR and DistgASR, and retrained the models using the dataset and codes officially released. The results, which provide a performance baseline, are documented in Table 8. To control for potential confounding effects of different network settings (e.g., the use of LeakyReLU), we restructured DistgSSR and DistgASR into template networks and integrated our PFE methodology, yielding DistgSSR-PFE and DistgASR-PFE. In the detailed implementation, PFE is deployed on the Disentangling block of the DistgSSR-PFE network and on both the Disentangling block and Disentangling group of the DistgASR-PFE network. The experimental findings, as outlined in Table 8, clearly demonstrate that our PFE can enhance network performance.
@@ 7 Conclusion
We have presented a novel adaptive probabilistic-based feature embedding for LFs. To verify its effectiveness, we evaluated its performance on compressive LF reconstruction and LF denoising, respectively. For compressive LF imaging, we proposed a physically interpretable cycle-consistent framework that incorporates probabilistic-based feature embedding. Our approach can reconstruct 4-D LFs from 2-D measurements with higher quality, improving the PSNR value by approximately 4.5 dB while preserving the LF parallax structure better than state-of-the-art methods. Furthermore, we demonstrate the efficacy of our probabilistic-based feature embedding approach for LF denoising through a carefully designed network. Extensive experiments on diverse datasets demonstrate that our method outperforms state-of-the-art approaches by up to $1.1$ dB.
In future work, the application of our proposed probabilistic-based feature embedding could be potentially broadened to include other tasks within the domain of LF image processing, such as depth estimation. Moreover, our proposed CR-Net exhibits significant potential for direct integration in reconstruction-based LF compression tasks Hou et al. (2018). The inherent capabilities of CR-Net could be leveraged to gather, compress, transmit, and subsequently reconstruct compressed sensing images. This process would yield high-quality light fields, all while necessitating less data.